OpenStudy (anonymous):

I need to know how to derive the zeros of the function f(x) = x^4-4x^3+30x^2-4x+29. I graphed it with a graphing calculator but it only showed as simple parabola, I need to find the complex zeros.

OpenStudy (anonymous):

there are no real zeros, that is why

OpenStudy (anonymous):

that is \[\{i, -i, 2+5i, 2-5i\}\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

perhaps you were supposed to guess i then if i works, you know that -i must be a zero as well. then you know that one factor is \[x^2+1\] so factor it out and get \[(x^2+1) (x^2-4 x+29)\] and solve the second one

OpenStudy (anonymous):

but how was the x^2-4x+29 obtained?

OpenStudy (anonymous):

factoring. once you know that i is a zero and also -i that means you have \[x^4-4x^3+30x^2-4x+29=(x^2+1)\times (\text{something)}\] so you just need to find the something. you can do it by factoring, or else by dividing

OpenStudy (anonymous):

you know "something" must look like \[x^2+ax+29\] because of the x^4 and the 29, and you know a = -4 because you have to end up with -4x and that is the only way to get it

OpenStudy (anonymous):

why do you have to end up with -4x?